# Moment Tensor Potential and Equivariant Tensor Network Potential with explicit dispersion interactions

### v1

@@ III.3.4 Molecular Dynamics
Table 8: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHarris and Trappeniers (1980).
To predict RDFs, the NpT simulation was performed for 150 ps at 92 K and 0.13 bar; subsequently, the ensemble was switched to NVT and the temperature was equilibrated for 10 ps, then the RDF was averaged over 100 ps in the NVE simulation. The resulting deviations of the RDFs from the experimentally observedOobatake, Hayashi, and Machida (1990) are reported in Table 9 and are nearly the same for all MLIPs, similar to the results obtained for CCl4.
MTP MTP+D2 ETN ETN+D2
MAPE(g) 0.036 0.027 0.035 0.031
$\pm$ 0.017 $\pm$ 0.012 $\pm$ 0.010 $\pm$ 0.011
MAPE(r) 0.019 0.019 0.019 0.025
$\pm$ 0.018 $\pm$ 0.011 $\pm$ 0.016 $\pm$ 0.015
Table 9: MAPE and one-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentOobatake, Hayashi, and Machida (1990).
Note that the pressure conditions of the simulations for density and RDF calculations differ significantly from those during the MD simulations in the AL process, yet the results are highly accurate, which indicates excellent transferability of the trained MLIPs.
III.4 Toluene modeling
As was demonstrated for liquid methane and carbon tetrachloride, MLIP(l) with extended $R_{\rm cut}$ can accurately capture dispersion interactions. However, methane and carbon tetrachloride molecules possess high symmetry, and the covalent and dispersion interactions are well separated in the interatomic distance scale. Now we proceed with the testing of MLIPs on liquid toluene, where the molecule is less symmetrical, the dispersion interactions are strongly dependent on the mutual orientation of molecules, are not clearly separated in the interatomic distance scale.
III.4.1 Datasets generation
All ab initio simulations for toluene were performed using FHI-aimsBlum et al. (2009) software with intermediate settings for basis set and integration grids for both atomic types and a 2$\times$2$\times$2 k-point grid. All simulations were carried out at the PBE-TS level of theory.
The AIMD simulation was performed in the NVT ensemble for 400 fs, at 298 K. The MD conditions in the AL were set to 298 K and 1 atm. At the end of the AL, the training set accumulated 520 configurations with sizes ranging from 15 to 18 Å, each consisting of 13 molecules.
III.4.2 Training and validation
Training and validation errors averaged over ensembles of MLIPs are shown in Figure 14. Judging by the energy validation errors, all the MLIP+D models are more accurate than their local counterparts, while this trend is not observed when looking at force errors.
Figure 14: Training and validation errors obtained for liquid toluene configurations.
III.4.3 Binding curves
The dimers proposed for this test are illustrated in Figure 15. They include dimers with and without stacking interactions, t-shaped dimers, and dimers with molecules lying in the same plane.
Figure 15: Dimers of toluene used for constructing binding curves and minima depths and positions. The labels denote mutual orientations of molecules: "P" – parallel, "T" – t-shaped, "IP" – in plane.
MAPE and RMSE, averaged over the ensembles of MLIP and dimers, are presented in Figure 16(a). The errors obtained are significantly higher than for other systems studied, including the modeling of CCl4 with MLIPs with small $R_{\rm{cut}}$. At the same time, there is a considerable improvement in minimum energy and FWHM accuracy for MLIP+D’s compared to MLIP(l)’s, especially for the ETN family of potentials.
Figure 16: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
The errors are strongly dependent on the mutual orientation of the molecules (Figure 16(b)), even to a greater extent than for CCl4. The highest MAPEs correspond to the most strongly interacting dimers with staking interactions: P-1, P-2, and P-3. In addition, IP-1, IP-2, and IP-3 also exhibit errors higher than average.
Figure 17: MBE for predicting binding curves minima depth (E), position (r) and full-width at half-minimum (FWHM), averaged over ensembles of potentials and dimers.
According to MBE (Figure 17), the ETN family of potentials tends to strongly overbind dimers, while MTP-based potentials show moderate underbinding. At the same time, the ETN-based potentials significantly underestimate the FWHM, while the MTP-based ones give more accurate predictions. Similarly to CCl4 and CH4, both MLIP(l)’s and MLIP+D’s accurately predict the position of the minima.
The large errors observed for the ETN family of potentials result from strong unphysical oscillations (Figure 18). This behavior is observed for all the MLIPs, but is more severe in the case of the ETN-based ones. Those oscillations are weaker for MLIP+D’s compared to MLIP(l)’s.
Figure 18: Binding curves of P-2, T-1 and IP-1 toluene dimers, predicted with PBE-TS (solid blue line) and with MLIPs (dashed line); the predicted binding curves are averaged over ensembles of MLIPs, and are demonstrated together with one-$\sigma$ confidence interval. All binding curves are plotted as functions of the distance between centers of mass (CM) of individual molecules.
Similarly to CCl4 and CH4, ETN+D3 does not outperform ETN+D2 in this test and, for this reason, will not be used in further simulations.
III.4.4 Molecular Dynamics
For MD simulations, a 30 Å simulation box consisting of 104 molecules was used. The density was equilibrated in the NpT ensemble for 200 ps with a 1 fs time step, at 298 K and 1 atm. The resulting density value was obtained by averaging over the last 100 ps and compared to the experimental valueHarris (2000) of 0.862 g/cm3. Figure 19(a) presents the density MAPE which is significantly smaller for MLIP+D’s than for MLIP(l)’s.
@@ III.4.4 Molecular Dynamics
The density as a function of time is illustrated in Figure 19(b) where the trajectories averaged over ensembles of MLIPs are shown together with the one-$\sigma$ confidence intervals. The density predicted by ETN exhibits a large uncertainty, indicating large deviations in the MLIP training results. Furthermore, the mean density evaluated with ETN significantly overestimates the experimental value. This is likely a consequence of the strong overbinding of molecules observed in the binding curve test (Figure 17). For the ETN + D2 potential, these issues are resolved, and a density value closer to the experiment is achieved together with a lower spread. MTP also shows higher error and uncertainty compared to MTP+D2, although the difference is not as dramatic as for ETN and ETN+D2.
Figure 19: a) Density MAPE and its one-$\sigma$ confidence interval, compared to experiment; b) density as a function of time together with one-$\sigma$ confidence interval; c) predicted RDFs (green and brown lines) and experimental RDFs (solid blue line); obtained for C7 (methyl) and C4 (opposite to the methyl group) carbons, averaged over ensemble of MLIPs, together with one-$\sigma$ confidence intervals.
@@
For RDF calculation, a 10 ps equilibration in the NVT ensemble was conducted for a simulation cell with the converged density from the previous simulation. After this, the ensemble was switched to NVE and 100 ps MD was performed with a 1 fs time step. During the NVE run, the configurations were collected every 50 steps and post-processed using MDAnalysisMichaud-Agrawal et al. (2011). The RDFs were calculated based on the C4 carbon atoms (opposite to the methyl group) and the C7 carbon atoms (methyl carbon). Three types of RDF were computed: C7–C4, C4–C4, and C7–C7. Note that in the C7-C4 RDF the intramolecular contribution was subtracted. The resulting RDFs were compared with the experimentFalkowska et al. (2016) and presented in Figure 19(c).
@@
RDFs obtained with all the MLIPs, except ETN, show qualitative agreement with the experiment. In contrast, the average C4–C4 RDF predicted with ETN has a false maximum. In addition, individual RDFs predicted by individual ETN potentials in the ensemble demonstrate significant spread from each other, leading to broad uncertainty regions. For the ETN+D2 potential, the width of the uncertainty region is substantially reduced and a better agreement with the experiment is achieved. The difference between MTP and MTP+D2 is less pronounced than for ETN and ETN+D2. However, a reduction in the uncertainty region is still observed, along with an improved alignment between the mean and experimental RDFs.
@@ IV Conclusions
We studied the impact of explicit incorporation of dispersion interactions in MTPs and ETN potentials via the D2 and D3 correction schemes. The impact was measured based on training and validation errors, errors in predicting binding curves, density, and RDFs of a liquid phase of three systems—carbon tetrachloride (CCl4), methane (CH4), and toluene.
@@
Another finding is that for the two relatively simple systems, CCl4 and CH4, the improvement in accuracy can instead be achieved by simply increasing the cutoff radius of the local MTP and ETN to $7.5\text{\AA}$, except for the full-width at the half-minimum metrics, in which the MLIP+D’s are still more accurate. We note that we did not systematically study the effect of increasing the cutoff radius beyond $7.5\text{\AA}$ (due to the fast increase in the cost of collecting the appropriate training set). For the toluene system, which has a sufficiently complex geometry, a simple increase in the cutoff radius was not enough to achieve the accuracy of the MLIP+D models, especially for the ETN model.

### v2

@@ III.4 Densities
Toluene 0.12 0.100 0.21 0.14 -
$\pm$ 0.02 $\pm$ 0.006 $\pm$ 0.06 $\pm$ 0.01
[…]
The two key findings during the modeling of the CCl4 system using all MLIPs with cutoff 5.5 Åare given below. First, the MTP potential does not maintain stable molecular dynamics in the liquid phase, as we observed the density collapsing to near zero values. However, this issue is resolved by explicit incorporation of dispersion interactions in MTP+D2, resulting in a 25.6% deviation from the experimentally measured density Hien, Ponter, and Peier (1978) of 1.594 g/cm3 (Figure 6 (a), Table 4). Unlike MTP, the ETN potential is capable of preserving the liquid phase, likely due to the overbinding of molecules, which was demonstrated in the dimer binding curves test (Figure 4 (b)). Additionally, ETN+D2 and ETN+D3 have a higher accuracy compared to ETN, achieving a deviation of approximately 22% from the experimental density. Secondly, unlike the results of the binding curves tests, ETN+D3 does not outperform ETN+D2 in predicting density.
When the cutoff radii were extended, we obtained a substantial improvement over the MLIPs with smaller cutoff radii, as can be seen from the density values and MAPEs (Figure 6 (b), Table 4). Furthermore, similarly to the binding curves, the errors of all MLIPs are close to each other. Thus, we demonstrated that it is sufficient to introduce an $R_{\rm cut}$ capturing the major part of long-range interactions for MLIP(l)’s for accurate modeling of liquid carbon tetrachloride, since the difference in accuracy between the MLIP(l) and MLIP+D potentials is minor for all properties calculated.
In view of all the results for CCl4, for the other two systems, we decided to limit the tests for MLIP+D’s with $R_{\rm cut}=6$ Å and MLIP(l)’s with $R_{\rm cut}=7.5$ Å.
@@
The density of the CH4 system was compared to the experimental value Harris and Trappeniers (1980) of 0.4256 g/cm3 (Figure 6 (c), Table 4). The errors obtained for all MLIPs are small and close in magnitude. Furthermore, no significant differences in the accuracy of MLIP(l)’s and MLIP+D’s are observed similarly to the modeling of CCl4. In contrast, the density of toluene predicted by ETN exhibits a large uncertainty, indicating substantial deviations in the MLIP training results (Figure 6 (d), Table 4). Furthermore, the mean density evaluated with ETN significantly overestimates the experimental value Harris (2000) of 0.862 g/cm3. For the ETN + D2 potential, these issues are resolved, and a density value closer to the experiment is achieved together with a lower spread. MTP also shows higher error and uncertainty compared to MTP+D2, although the difference is not as dramatic as for ETN and ETN+D2.
To sum up, we observed that the MLIP+D’s were more accurate for density calculations than MLIP(l)’s for CCl4 when the cutoff radii were set to 5.5 Å. However, when the cutoff radii of MLIP(l)’s were extended, their accuracy in modeling CCl4 and CH4 improved to the level of MLIP+D’s and even surpassed MTP+D2 for MTP potentials. In contrast, for toluene, ETN+D2 significantly outperformed ETN even when its cutoff radius was extended. This is likely attributed to the unphysical oscillations observed in toluene binding curves (Figure 5), which were significantly reduced after incorporating the dispersion correction.
III.5 Radial distribution functions
Finally, we calculated RDFs, which were obtained by averaging over 100 ps of the NVE MD trajectories after the equilibration of the density in the NpT ensemble and the temperature in the NVT ensemble. The details of the simulations can be found in Section III.1. The RDFs calculated with MLIPs are shown in Figure 7 and Table 5 contains the MAPEs for extrema positions and intensities, compared to those obtained experimentally for CCl4 Lowden and Chandler (1974), CH4 Oobatake, Hayashi, and Machida (1990), and toluene Falkowska et al. (2016).
Figure 7: Radial distribution functions obtained from experiment Lowden and Chandler (1974); Oobatake, Hayashi, and Machida (1990); Falkowska et al. (2016) and with MLIPs for: a) CCl4 using MLIPs with cutoff 5.5 Å; b) CCl4 using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å; c) CH4 using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å; d) toluene using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å.
Error MTP MTP+D2 ETN ETN+D2 ETN+D3
[…]
$\pm$ 0.03 $\pm$ 0.04 $\pm$ 0.05 $\pm$ 0.05
MAPE(r) - 0.04 0.03 0.03 0.03
$\pm$ 0.02 $\pm$ 0.02 $\pm$ 0.02 $\pm$ 0.02
[…]
$\pm$ 0.08 $\pm$ 0.05 $\pm$ 0.09 $\pm$ 0.09
MAPE(r) 0.03 0.03 0.03 0.03 -
$\pm$ 0.01 $\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.01
CH4 MAPE(g) 0.04 0.03 0.04 0.03 -
$\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.01 $\pm$ 0.01
MAPE(r) 0.02 0.02 0.02 0.03 -
$\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.02 $\pm$ 0.02
[…]
All MLIPs with $R_{\rm cut}=5.5$ Å give nearly equivalent RDFs for CCl4 (Figure 7 (a), Table 5), at the same time exhibiting a shift compared to the experimental data due to an underestimation of the density. In contrast to the results obtained for the density, we do not observe a substantial improvement in accuracy after incorporating dispersion corrections in the present case. Note that the RDF was not calculated with MTP as it failed to maintain the liquid phase in the MD simulations. When CCl4 is modeled with MLIP(l)’s with a cutoff of 7.5 Å, and MLIP+D’s with cutoff 6 Å, the predictions of all MLIPs are close to each other (Figure 7 (b), Table 5), similarly to the calculations using MLIPs with smaller cutoffs.
For the CH4 system, similarly to the accurate density predictions, the resulting RDFs are close to the experimental ones and are nearly the same for all (see Figure 7 (c), Table 5). Additionally, we note that the pressure conditions of the simulations for density and RDF calculations differ significantly from those during the MD simulations in the AL process, yet the results are highly accurate, indicating excellent transferability of the trained MLIPs.
@@ III.5 Radial distribution functions
For modeling toluene, the configurations were collected every 50 fs and post-processed using MDAnalysis Michaud-Agrawal et al. (2011). The RDFs were calculated based on the C4 carbon atoms (opposite to the methyl group) and the C7 carbon atoms (methyl carbon). Three types of RDFs were computed: C7–C4, C4–C4, and C7–C7 (Figure 7 (d)). Note that in the C7–C4 RDF the intramolecular contribution was subtracted.
@@
RDFs obtained with all the MLIPs, except for ETN, show qualitative agreement with the experiment. In contrast, the average C4–C4 RDF predicted with ETN has a false maximum. In addition, individual RDFs predicted by individual ETN potentials in the ensemble demonstrate significant spread from each other, leading to broad uncertainty regions. For the ETN+D2 potential, the width of the uncertainty region is substantially reduced and a better agreement with the experiment is achieved. The difference between MTP and MTP+D2 is less pronounced than for ETN and ETN+D2. However, a reduction in the uncertainty region is still observed, along with an improved alignment between the mean and experimental RDFs. As in the case of density prediction, we associate the higher accuracy of the MLIP+D’s with reduced oscillations observed on the binding curves (Figure 5).
Overall, we observed that the CCl4 and CH4 RDFs were not sensitive to the incorporation of dispersion corrections in the functional form of the MLIPs. However, it had a significant effect on the RDF for toluene, highlighting that this effect depends on system complexity.
Additionally, we want to emphasize that the MLIP+D models still have advantages for the extended-cutoff MLIP(l)’s. In fact, employing a smaller $R_{\rm cut}$ while incorporating explicit dispersion corrections improves the computational efficiency of potential fitting and subsequent simulations. This efficiency stems from the reduced computational cost of DFT calculations for smaller configurations, as well as the faster training and evaluation of energies and forces with potentials that have a smaller cutoff radius.
It should be also noted that we started this study by calculating binding energies, density, and RDF using MLIPs with $R_{\rm cut}=5.5$ Å for the CCl4 system. We found that ETN+D3 outperformed ETN+D2 only for the CCl4 binding curves and, for this reason, we did not use ETN+D3 for other tests provided here.
Finally, the remaining deviations between the predicted densities and RDFs and experimental ones arises not only from MLIP fitting errors but also from intrinsic limitations of the PBE density functional, the dispersion correction schemes, and the neglect of many-body and quantum nuclear effects. While addressing these factors could potentially enhance the accuracy compared to experiment, it was beyond the scope of the present study.
@@ IV Conclusions
We studied the impact of explicit incorporation of dispersion interactions in MTPs and ETN potentials via the D2 and D3 correction schemes. The impact was measured based on training and validation errors, errors in predicting binding curves, density, and RDFs of a liquid phase of three systems – carbon tetrachloride (CCl4), methane (CH4), and toluene.
@@
Another finding is that for the two relatively simple systems, CCl4 and CH4, the improvement in accuracy can instead be achieved by simply increasing the cutoff radius of the local MTP and ETN to 7.5 Å, except for the full-width at the half-minimum metrics in CH4, in which the MLIP+D’s are still more accurate. We note that we did not systematically study the effect of increasing the cutoff radius beyond 7.5 Å (due to the fast increase in the cost of collecting the appropriate training set). For the toluene system, which has a sufficiently complex geometry, a simple increase in the cutoff radius was not enough to achieve the accuracy of the MLIP+D models, especially for the ETN model.
Supplementary material
See the supplementary material for additional data and supporting discussion of statements made in the main text.